OpenStudy (anonymous):

Translate this problem to an equation and solve. The area of a rug is 36 square feet. The length of the rug is 5 feet longer than the width. What is the width of the rug?@johnweldon1993

OpenStudy (anonymous):

the area is length*breadth length=5+breadth so area =breadth*(breadth+5) 36=b^2+5b so solving it we get b=4

OpenStudy (anonymous):

so length =b+5=4+5=9

OpenStudy (anonymous):

but they want feet is the answer 9 feet @johnweldon1993

OpenStudy (anonymous):

yes one root is -9 but it cant be true as breadth can never be negetive

OpenStudy (anonymous):

thta not the answer aditya96

OpenStudy (johnweldon1993):

It trys to trick you....you went and solved for Length...it asks for Width :) @aditya96 did it correctly...but you gave the LENGTH as the answer...the answer is 4 as @aditya96 has shown above

OpenStudy (anonymous):

but @shortie1515 says breadth is 9

OpenStudy (anonymous):

@johnweldon1993 so its 4 feet right

OpenStudy (johnweldon1993):

No, You have done the correct procedure @aditya96 Length = 9 feet Width = 4 feet You have solved for both correctly...but the answer @shortie1515 should have taken...is 4 feet

OpenStudy (johnweldon1993):

Yes @shortie1515 the answer is 4 feet